OpenStudy (anonymous):

A coin container has 10 more nickels than dimes and together the nickels and dimes are worth $3.50. How any nickels and dimes are in the container?

OpenStudy (anonymous):

let n be the number of nickels and d the number of dimes n = d + 10 by looking at the first sentence 0.1d + 0.05n = 3.50 since nickels are 5c ea and dimes 10c a piece. Do you like elimination or substitution better?

Directrix (directrix):

Hitchhiking on bluepig's work: n = d + 10 0.1d + 0.05n = 3.50 --> Multiply by 100 -> 10d + 5 n = 350 n = d + 10 10d + 5n = 350 10d + 5 (d+10) = 350 10d + 5d + 50 = 350 15d + 50 = 350 15d = 300 d = 20 ---> 20 dimes n = d + 10 n = 20 + 10 = 30 ---> nickels